OpenStudy (boxman61):

let logb2=1.5 and logb8=0.25 find logb16 need to know process please..

OpenStudy (luffingsails):

logb(M*N) = logb(M) + logb(N) logb(16) = logb(2*8) = logb(2) + logb(8) = 1.5 + 0.25 = 1.75

OpenStudy (boxman61):

just like that its easy... thank you

OpenStudy (luffingsails):

Glad to help.